# DriveSOTIF: Advancing SOTIF Through Multimodal Large Language Models

### v2

@@ Abstract
Human drivers possess spatial and causal intelligence, enabling them to perceive driving scenarios, anticipate hazards, and react to dynamic environments. In contrast, autonomous vehicles lack these abilities, making it challenging to manage perception-related Safety of the Intended Functionality (SOTIF) risks, especially under complex or unpredictable driving conditions. To address this gap, we propose fine-tuning multimodal large language models (MLLMs) on a customized dataset specifically designed to capture perception-related SOTIF scenarios. Benchmarking results show that fine-tuned MLLMs achieve an 11.8% improvement in close-ended VQA accuracy and a 12.0% increase in open-ended VQA scores compared to baseline models, while maintaining real-time performance with a 0.59-second average inference time per image. We validate our approach through real-world case studies in Canada and China, where fine-tuned models correctly identify safety risks that challenge even experienced human drivers. This work represents the first application of domain-specific MLLM fine-tuning for SOTIF domain in autonomous driving. The dataset and related resources are available at github.com/s95huang/DriveSOTIF.git
@@ I Introduction
Perception systems in autonomous vehicles (AVs),which are responsible for detecting, classifying, and predicting based on environmental inputs, are especially vulnerable to SOTIF-related risks. Factors such as sensor limitations in adverse weather, difficulty recognizing unusual road users, or rare and complex scenarios can result in unsafe decisions [2]. The dynamic, unpredictable nature of real-world driving further complicates matters: sudden obstacles or erratic behavior by other road users can easily overwhelm current algorithms [3].
Identifying and mitigating SOTIF risks is therefore essential for safe AV operation. Current methods primarily use rule-based approaches to quantify perception-related risks, for example by measuring output entropy and classifying risks into predefined categories [4]. Learning-based methods, on the other hand, suffer from black-box behavior, poor generalization to edge cases, and vulnerability to adversarial conditions. In contrast, human drivers intuitively assess risk and uncertainty for specific objects as they drive, drawing on personal experience, knowledge, and common sense—capabilities that remain absent from existing automated systems.
Recent advances in multimodal large language models (MLLMs) provide new tools to address these cognitive gaps. MLLMs offer three distinct advantages over traditional approaches: (1) open-world cognition that leverages pre-trained knowledge to interpret novel scenarios; (2) causal reasoning that traces how environmental factors cascade into safety risks; and (3) contextual inference that interprets social dynamics and behavioral patterns beyond conventional algorithmic reach [5, 6]. For example, an MLLM can infer that sun glare may obscure lane markings, recognize that people gathering with umbrellas likely indicates a bus stop, or understand how rain, construction, and nighttime conditions interact to create heightened safety risks. Crucially, MLLMs provide explainable risk reasoning via natural language, articulating why scenarios pose specific risks and recommending actionable responses—capabilities essential for safety validation and user trust.
@@ II-A Perception-Related SOTIF
Perception-related SOTIF risks occur when an autonomous vehicle’s perception system functions correctly at both the hardware and software levels yet fails to accurately interpret the environment due to external factors [3]. Because perception underpins downstream tasks such as decision-making and planning, errors at this stage can propagate, leading to hazardous outcomes including false emergency braking, unsafe lane changes, or collisions. These challenges are particularly acute in adverse conditions—rain, snow, fog, glare—or in the presence of rare or unusual objects [7], especially when vehicles operate outside their defined Operational Design Domain (ODD) [11, 12].
Mitigation approaches can be broadly categorized as rule-based or learning-based. Rule-based methods—such as Fault Tree Analysis (FTA) [4], System-Theoretic Process Analysis (STPA) [13, 14], and formal verification [15]—use structured frameworks to model how specific failures can lead to unsafe behaviors. While effective for high-level system modeling, these methods often fail to capture the unpredictable and dynamic aspects of real-world environments [16].
Learning-based and probabilistic techniques focus on modeling perception uncertainty using data-driven methods. Bayesian Risk Decomposition [17], deep ensembles [7, 3, 18], and sensor fusion have shown promise in quantifying and mitigating uncertainty in perception outputs [4]. However, these models typically function as black boxes, limiting interpretability and making it difficult to diagnose failure modes. Their performance also tends to degrade in rare or edge-case scenarios that fall outside typical training distributions.
II-B MLLMs in Autonomous Driving and Dataset Challenges
Multimodal large language models (MLLMs) are attracting increasing attention in autonomous driving due to their ability to jointly process visual and textual information [19]. This enables the integration of visual inputs—such as camera feeds—with map data, vehicle state, and textual cues, leading to a more holistic understanding of the driving environment [20, 21].
From a SOTIF perspective, MLLMs are especially promising for addressing perception limitations. Their cross-modal reasoning allows them to compensate when one sensor stream is degraded—for example, when camera images are obscured by glare or LiDAR signals are weakened by fog. By integrating complementary sensory inputs, MLLMs can infer scene context, improving situational awareness and reducing the risk of misinterpretation under challenging conditions [22, 21, 23].
@@ II-B MLLMs in Autonomous Driving and Dataset Challenges
However, the effective deployment of MLLMs in SOTIF applications is currently hampered by dataset limitations. Most available SOTIF-related datasets, summarized in Table I, are focused on object detection tasks. For example, PeSOTIF [7] and CODA [8] target 2D detection in rare and long-tail scenarios, while SOTIF-PCOD [9] provides synthetic LiDAR point clouds. Of these, only CODA-LM [10] includes textual annotations for vision-language models, but its primary focus is not direct SOTIF risk assessment.
In addition, existing MLLM-focused driving datasets present further challenges. Datasets like DRAMA [24] and NuScenes-QA [25], while large and well-annotated, are geographically or scenically constrained, which limits their ability to generalize to new environments. Many also rely on template-based question-answer generation [25, 26], resulting in interactions that lack the nuance and diversity of natural human queries. This restricts the models’ capacity for open-ended reasoning and complex situational judgment, which are essential for safety-critical autonomous driving applications.
@@ III DriveSOTIF
We introduce DriveSOTIF, a domain-specific dataset designed to capture perception-related SOTIF scenarios for autonomous driving. As illustrated in Figure 1, our approach combines dataset generation with supervised fine-tuning of pre-trained MLLMs.
@@ III-A Image Selection
For this research, we utilize the first batch of images from PeSOTIF [7], an object-detection dataset created to study the perception SOTIF problem in long-tail traffic scenarios. Specifically, the PeSOTIF dataset contains natural and artificial scenarios of perception SOTIF scenarios caused by environmental factors such as adverse weather conditions (e.g., rain, snow, fog), challenging lighting conditions (e.g., direct sunlight, glare from headlights), and object-related factors like unusual object appearances or unexpected object behaviors.
Today’s passenger vehicles often contain multiple cameras, posing a challenge for VQA tools that need to incorporate these inputs to effectively assist in BEV-based perception SOTIF problems. For a dataset newly introduced to SOTIF, we begin with simple image-question pairs. This approach allows for comparisons with existing VQA datasets and algorithms and encourages further research in this area.
@@
Open-ended Q&A target higher-level reasoning, including causal inference and scenario explanation. These questions do not follow a fixed structure and can have multiple correct answers, offering more flexibility and depth in responses. These questions are focused on SOTIF risk identification, explanation, and providing recommended actions.
To ensure a balanced representation of question types, we include five questions per image, with 2–3 closed-ended and 2–3 open-ended. This approach ensures a diverse set of questions and promotes a comprehensive evaluation of the dataset.
@@
To automatically generate the VQA and image captioning datasets, we leverage a multi-agent framework using vision-capable large language models (LLMs), specifically OpenAI’s GPT-4v, GPT-4o, and Anthropic’s Claude 3 Opus. The dataset generation pipeline, illustrated in Figure 2, includes distinct roles for image captioning, question generation, answer generation, and validation/post-processing. All prompts used in the dataset generation and validation process are provided in S.III of the supplemental material.
@@
Answer Generation Agent: In the VQA dataset format, each question is answered by ten responses from ten annotators. While this approach ensures accuracy, it is both costly and time-consuming. To streamline this process, the answer generation agent is designed to provide ten answers for each question, given an image. To achieve this, we employ a model aggregation method that utilizes GPT-4o and Claude3-Opus. For each image, the image file and associated questions are fed into the LLMs. Instead of sequentially querying a single model multiple times, the system generates five answers for each question in a single pass through each model. This method is more efficient and cost-effective compared to sending the questions individually. Similarly, the answer generation agent outputs metadata for each answer, including the confidence score and answer type.
Validation and Post-processing Agent: As shown in Figure 2, newly generated answers are then sent to the validation and post-processing agent, which contains three post-processing mechanisms. The first and second mechanisms are to ensure that captions and questions generated are related to the image and perception SOTIF context. The third mechanism is to ensure that answers generated are correct based on the input image and question. If any inconsistencies are found, the agent regenerates the captions, questions or answers as needed. This validation process leverages OpenAI’s GPT4.1 as the LLM backend to ensure high-quality outputs that align with the perception SOTIF context.
@@ III-C Dataset Overview
The multi-agent system employed varying temperature settings during generation to ensure diversity across questions, answers, and captions. Following automated generation, all data entries underwent structural validation and format standardization, with close-ended and open-ended questions automatically categorized using metadata from the generation process. A human review of 595 samples, selected using finite population sampling [^2] More details on the finite population sampling method can be found in S.IV of the supplemental material.., indicated a 3-4% error rate, which is acceptable for automated pipelines. An additional round of manual review addressed edge cases and subtle inconsistencies that automated systems might miss or generate.
The final DriveSOTIF dataset contains 1,114 images, 1,114 captions, and 5,570 question-answer pairs  [^3] Collecting SOTIF-related scenarios is challenging due to their rarity and unpredictability in real-world settings. We are actively collaborating with PeSOTIF authors to expand the dataset with new scenarios and modalities.. The dataset adopts COCO Caption format [27] for image descriptions and VQAv2 format [28] for question-answering tasks. We partition the data using a 90/10 train-test split, as detailed in Table IV.
Table III demonstrates the dataset’s strong alignment with perception-related SOTIF challenges. Quantitatively, caption length averages 18-19 words compared to 10.5 words in COCO Caption, providing substantially richer contextual information essential for understanding complex driving scenarios. Question difficulty maintains excellent balance across easy (39.9%), medium (40.2%), and hard (19.9%) categories. Similarly, question formats split evenly between close-ended (45.7%) and open-ended (54.3%) types, which facilitates assessment of both factual accuracy and reasoning capabilities. Answer types span analysis (36.1%), yes/no and multiple-choice (28.1%), recommendations (17.8%), counting (17.0%), and identification (1.0%) tasks. This diversity enables comprehensive benchmarking of models’ ability to handle both fundamental perception tasks and higher-level causal reasoning and safety recommendations specific to SOTIF.
@@ IV-A Benchmarking
For image captioning benchmarking, we evaluate the Blip [33] and Blip2 models [34] (pre-trained on the COCO Caption dataset [27]) across various model sizes. For VQA tasks, we benchmark a total of Four open-sourced models, including LLaVA 1.5 [35], Qwen2/2.5-VL [36, 37], and InterVL3 [38].
@@ IV-B Model Fine-tuning
To enhance the performance of our benchmarked models, we adopt Parameter-Efficient Fine-Tuning (PEFT) techniques such as Low-Rank Adaptation (LoRA) [39] and Quantized LoRA (QLoRA) [40]. LoRA introduces low-rank matrices into each Transformer layer, allowing the model to adapt with far fewer trainable parameters while keeping the original model weights frozen. QLoRA builds on this by incorporating quantization, which further reduces memory requirements and enhances efficiency, all while maintaining performance levels. This makes it easier to adapt LLMs for specialized tasks without needing extensive hardware resources.
For image captioning, we fine-tuned Blip and Blip2 models using the LAVIS framework [41]. For VQA, we utilized LLaVA [35] and LLaMA-Factory [42] frameworks to fine-tune LLaVA 1.5, Qwen2/2.5-VL, and other leading vision-language models. Hyperparameters, random seeds, and training configurations were kept consistent across experiments for fair comparison (details in S.II of the supplemental material).
@@
Method Size BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr SPICE
Base FT Base FT Base FT Base FT Base FT Base FT Base FT Base FT
Blip Small 0.348 0.659 0.219 0.464 0.132 0.318 0.082 0.217 0.134 0.208 0.315 0.414 0.138 0.411 0.086 0.143
Large 0.349 0.671 0.227 0.463 0.139 0.312 0.083 0.206 0.137 0.210 0.324 0.399 0.140 0.410 0.091 0.145
@@ IV-C1 Qualitative Analysis
For VQA, baseline models typically succeed on simple, factual questions but often fail to provide detailed, situation-specific reasoning. Fine-tuned models demonstrate better understanding of complex queries. For example, when prompted about rain-related perception issues, the fine-tuned model highlights specific challenges like reflections and distortions, offering a deeper analysis of SOTIF risks. In unusual scenarios—such as an overturned vehicle—fine-tuned models identify relevant perception challenges (e.g., blocked sightlines or misclassification risks) that baseline models may overlook.

### v3

@@ Abstract
Human drivers possess spatial and causal intelligence, enabling them to perceive driving scenarios, anticipate hazards, and react to dynamic environments. In contrast, autonomous vehicles lack these abilities, making it challenging to manage perception-related Safety of the Intended Functionality (SOTIF) risks, especially under complex or unpredictable driving conditions. To address this gap, we propose fine-tuning multimodal large language models (MLLMs) on a customized dataset specifically designed to capture perception-related SOTIF scenarios. Benchmarking results show that fine-tuned MLLMs achieve an 11.8% improvement in close-ended VQA accuracy and a 12.0% increase in open-ended VQA scores compared to baseline models, while maintaining real-time performance with a 0.59-second average inference time per image. We validate our approach through real-world case studies in Canada and China, where fine-tuned models correctly identify safety risks that challenge even experienced human drivers. This work represents the first application of domain-specific MLLM fine-tuning for the SOTIF domain in autonomous driving. The dataset and related resources are available at github.com/s95huang/DriveSOTIF.git
@@
©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
Perception systems in autonomous vehicles (AVs), which are tasked with detecting, classifying, and predicting based on environmental stimuli, are particularly vulnerable to SOTIF-related challenges. Issues such as sensor limitations in adverse weather, difficulty in recognizing unusual road users, or interpreting rare and complex scenarios can lead to unsafe decisions [2]. The unpredictable nature of real-world driving, such as sudden obstacles or erratic road users, can easily overwhelm current algorithms [3].
Detecting and addressing SOTIF risks are therefore essential for the safe operation of AVs. Existing methods primarily adopt rule-based approaches to quantify perception-related risks by measuring entropy levels and classifying them into specific categories [3, 4]. Learning-based methods, on the other hand, suffer from black-box behavior, poor generalization to edge cases, and vulnerability to adversarial conditions. In contrast, human drivers can naturally assess the risk and uncertainty with certain objects when driving. This capability of assessing SOTIF-related risk and selecting the best action based on personal experience, knowledge, common sense remains absent from existing literature.
Recent advances in multimodal large language models (MLLMs) offer new approaches for addressing these cognitive gaps. MLLMs provide three unique capabilities [5, 6]:
  • Open-world generalization: leveraging pre-trained knowledge to interpret unseen or novel scenarios.
  • Causal reasoning: tracing cause–effect relationships to understand how individual factors, or their combinations, contribute to specific outcomes.
  • Contextual understanding: interpreting meaning within specific contexts, such as social interactions and human behavioral patterns.
For example, an MLLM may infer that sun glare obscures lane markings or that rain, road construction, and nighttime driving jointly increase safety risks. Importantly, MLLMs can assess such scenarios and provide human-readable explanations of the hazards involved, along with recommended actions.
@@ II-A Perception-Related SOTIF
Perception SOTIF risks arise when an autonomous vehicle’s perception system functions correctly at the hardware and software levels but still fails to accurately interpret the driving environment due to external factors [3]. Since perception is fundamental for downstream decision-making and planning, any misinterpretation can propagate through the system, leading to unintended emergency braking, unsafe lane changes, or even collisions. These risks are further amplified by factors like sensor degradation due to rain, fog, snow, and glare, and uncommon object appearances [4]. Identifying and mitigating perception-related SOTIF risks is crucial for maintaining vehicle safety, particularly when an autonomous vehicle operates outside its predefined Operational Design Domain (ODD) [10, 11].
Current approaches to detect and mitigate perception-related SOTIF risks can be classified into rule-based and learning-based methods. Rule-based approaches, such as Fault Tree Analysis (FTA) [12], System Theoretic Process Analysis (STPA) [13, 14], and formal verification [15], attempt to identify perception-triggering conditions and system failures through structured models [14]. However, these methods struggle with dynamic and unpredictable environments, limiting their applicability to real-world driving conditions [16].
In contrast, learning-based and probabilistic uncertainty quantification techniques aim to detect and mitigate perception-related SOTIF risks through data-driven models. Methods such as Bayesian Risk Decomposition [17], deep ensemble [4, 3, 18], and sensor fusion techniques have been explored to quantify uncertainty in perception outputs and improve reliability [12]. Despite these advancements, learning-based methods often suffer from black-box behavior, making it challenging to interpret the underlying process, and poor generalization to edge cases, particularly in long-tail traffic scenarios.
II-B MLLMs for SOTIF Applications
Multimodal large language models (MLLMs) integrate visual feature encoders with text-based LLMs to enable cross-modal understanding and reasoning[19, 5, 6]. By combining image and textual information from other modules, such as vehicle states, safety, decision-making, planning, and control, the system can make more informed responses and better understand and adapt to dynamic driving scenarios [20, 21, 22]. These capabilities make MLLMs suitable for perception-SOTIF scenarios in unpredictable driving conditions [23, 24, 25, 26, 27, 21].
@@ II-B MLLMs for SOTIF Applications
Despite this potential, SOTIF-specific MLLM development remains limited by the datasets currently available. As shown in Table I, existing SOTIF datasets primarily focus on object detection tasks. In detail, PeSOTIF [4] and CODA [7] are designed for 2D detection in rare and long-tail scenarios, while SOTIF-PCOD [8] contains synthetic LiDAR point clouds. CODA-LM [9] includes text annotations for perception and driving suggestions but does not address SOTIF risk directly.
Without SOTIF-specific MLLM datasets, researchers must use general-purpose datasets with limited relevance to SOTIF challenges. Datasets like DRAMA [28] and NuScenes-QA [24] are large and well-annotated but are geographically constrained, making it difficult to generalize to new scenarios. Other datasets [24, 29] often use template-based question-answer generation, resulting in artificial interactions that don’t reflect natural human queries. As a result, models trained/fine-tuned on these datasets struggle with open-ended reasoning and complex situational judgment.
@@ III DriveSOTIF
Based on the literature review, it is clear that existing datasets fall short of supporting SOTIF-focused MLLM development. Most current resources are either centered on object detection or limited by geographic scope and template-driven annotations, which prevents them from capturing the diversity and unpredictability of real-world safety scenarios. To address this gap, we introduce DriveSOTIF, a domain-specific dataset designed to capture perception-related SOTIF scenarios for autonomous driving. As illustrated in Figure 1, our approach combines dataset generation with supervised fine-tuning of pre-trained MLLMs.
@@ III-A Image Selection
For this research, we utilize the first batch of images from PeSOTIF [4], an object-detection dataset created to study the perception SOTIF problem in long-tail traffic scenarios. Specifically, the PeSOTIF dataset contains natural and artificial scenarios of perception SOTIF challenges caused by environmental factors such as adverse weather conditions (e.g., rain, snow, fog), challenging lighting conditions (e.g., direct sunlight, glare from headlights), and object-related factors like unusual object appearances or unexpected object behaviors.
Today’s passenger vehicles are typically equipped with multiple cameras, which requires VQA models to fuse information from different viewpoints for BEV-based perception SOTIF tasks. For a dataset newly introduced to SOTIF, we begin with simple image-question pairs. This approach allows for comparisons with existing VQA datasets and algorithms and encourages further research in this area.
@@
Open-ended Q&A targets higher-level reasoning, including causal inference and scenario explanation. These questions do not follow a fixed structure and can have multiple correct answers, offering more flexibility and depth in responses. These questions are focused on SOTIF risk identification, explanation, and providing recommended actions.
To maintain balance across question types, we assign five questions to each image, with 2–3 closed-ended and 2–3 open-ended. This design ensures balanced question diversity for comprehensive evaluation.
@@
To automatically generate the VQA and image captioning datasets, we develop a multi-agent framework using vision-capable large language models (LLMs), specifically OpenAI’s GPT-4v, GPT-4o, and Anthropic’s Claude 3 Opus. The dataset generation pipeline, illustrated in Figure 2, includes distinct roles for image captioning, question generation, answer generation, and validation/post-processing. All prompts used in the dataset generation and validation process are provided in S.III of the supplemental material.
@@
Answer Generation Agent: In the VQA dataset format, each question is answered by ten responses from ten annotators. While this approach ensures accuracy, it is both costly and time-consuming. To streamline this process, the answer generation agent is designed to provide ten answers for each question, given an image. To achieve this, we employ a model aggregation method that utilizes GPT-4o and Claude3 Opus. For each image, the image file and associated questions are fed into the LLMs. Instead of sequentially querying a single model multiple times, the system generates five answers for each question in a single pass through each model. This method is more efficient and cost-effective compared to sending the questions individually. Similarly, the answer generation agent outputs metadata for each answer, including the confidence score and answer type.
Validation and Post-processing Agent: As shown in Figure 2, newly generated answers are then sent to the validation and post-processing agent, which contains three post-processing mechanisms. The first and second mechanisms are to ensure that captions and questions generated are related to the image and perception SOTIF context. The third mechanism is to ensure that answers generated are correct based on the input image and question. If any inconsistencies are found, the agent regenerates the captions, questions or answers as needed. This validation process leverages OpenAI’s GPT-4.1 as the LLM backend to ensure high-quality outputs that align with the perception SOTIF context.
@@ III-C Dataset Overview
The multi-agent system employed varying temperature settings during generation to ensure diversity across questions, answers, and captions. Following automated generation, all data entries underwent structural validation and format standardization.
with close-ended and open-ended questions automatically categorized using metadata from the generation process. A human review of 595 samples, selected using finite population sampling [^2] More details on the finite population sampling method can be found in S.IV of the supplemental material. indicated a 3-4% error rate, which is acceptable for automated pipelines. An additional round of manual review addressed edge cases and subtle inconsistencies that automated systems might miss or generate.
The final DriveSOTIF dataset contains 1,114 images, 1,114 captions, and 5,570 question-answer pairs  [^3] Collecting SOTIF-related scenarios is challenging due to their rarity and unpredictability in real-world settings. We are actively collaborating with PeSOTIF authors to expand the dataset with new scenarios and modalities.. The dataset adopts the COCO Caption format [30] for image descriptions and the VQAv2 format [31] for question-answering tasks. We partition the data using a 90/10 train-test split, as detailed in Table IV.
Table III demonstrates the dataset’s strong alignment with perception-related SOTIF challenges. Quantitatively, caption length averages 18-19 words compared to 10.5 words in COCO Caption, providing substantially richer contextual information essential for understanding complex driving scenarios. Question difficulty maintains excellent balance across easy (39.9%), medium (40.2%), and hard (19.9%) categories. Similarly, question formats split evenly between close-ended (45.7%) and open-ended (54.3%) types, which facilitates assessment of both factual accuracy and reasoning capabilities.
Answer types span analysis (36.1%), yes/no and multiple-choice (28.1%), recommendations (17.8%), counting (17.0%), and identification (1.0%) tasks. This diversity enables comprehensive benchmarking of models’ ability to handle both fundamental perception tasks and higher-level causal reasoning and safety recommendations specific to SOTIF.
@@ IV-A Benchmarking
For image captioning benchmarking, we evaluate the BLIP [36] and BLIP-2 models [37] (pre-trained on the COCO Caption dataset [30]) across various model sizes. For VQA tasks, we benchmark a total of four open-sourced models, including LLaVA 1.5 [38], Qwen2/2.5-VL [39, 40], and InterVL3 [41].
@@ IV-B Model Fine-tuning
To enhance model performance, we apply Parameter-Efficient Fine-Tuning (PEFT) using Low-Rank Adaptation (LoRA) [42]. LoRA introduces low-rank matrices into each Transformer layer, allowing the model to adapt with far fewer trainable parameters while keeping the original model weights frozen. This approach allows LLMs to be specialized for downstream tasks with substantially lower computational and hardware demands than full fine-tuning.
For image captioning, we fine-tuned BLIP and BLIP-2 models using the LAVIS framework [43]. For VQA, we utilized LLaVA [38] and LLaMA-Factory [44] frameworks to fine-tune LLaVA 1.5, Qwen2/2.5-VL, and other leading vision-language models. Hyperparameters, random seeds, and training configurations were kept consistent across experiments for fair comparison (details in S.II of the supplemental material).
@@
Method Size BLEU-4 METEOR ROUGE-L CIDEr SPICE
Base FT Base FT Base FT Base FT Base FT
BLIP Small 0.082 0.217 0.134 0.208 0.315 0.414 0.138 0.411 0.086 0.143
Large 0.083 0.206 0.137 0.210 0.324 0.399 0.140 0.410 0.091 0.145
@@ IV-C1 Qualitative Analysis
For VQA, baseline models succeed on close-ended questions but often fail to provide detailed, situation-specific reasoning. Fine-tuned models demonstrate better understanding of complex queries. For example, when prompted about rain-related perception issues, the fine-tuned model points out specific SOTIF risk factors like reflections and distortions caused by rain that can degrade sensor performance. Similarly, in rare scenarios like an overturned vehicle, the fine-tuned model recognizes that the unusual positioning blocks the driver’s field of view and may cause misclassification. In contrast, the baseline model response lacks such depth.